OpenStudy (xapproachesinfinity):

interesting integrals \[\int \sqrt{\frac{x}{1-x^3}}dx\] \[\int \frac{1}{\ln x}+\ln(\ln x) dx\]

OpenStudy (xapproachesinfinity):

Second I solved it! first still beats me

OpenStudy (serenity74):

integrals = cringes

OpenStudy (xapproachesinfinity):

any hint for that first one @imqwerty

imqwerty (imqwerty):

m thinking :) i havn't learned integrals in depth yet in institute tho

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int \frac{1}{\ln x}+\ln (\ln x)~dx }\) z=ln(x) dz = dx/x --> x dz = dx --> e^z du =dx \(\color{#000000 }{ \displaystyle \int \left(\frac{e^z}{z}+e^z\ln z\right)~dz }\) \(\color{#000000 }{ \displaystyle \int \left(\sum_{k=0}^{\infty}\frac{z^{k-1}}{k!}\right)dx+\int\left(e^z\ln z\right)~dz }\) the first is series, ez. second probably integ. by parts

OpenStudy (solomonzelman):

nope, same thing, \(\color{#000000 }{ \displaystyle \int\left(e^z\ln z\right)~dz =e^z \ln z-\int\frac{e^z}{z}dz}\)

OpenStudy (solomonzelman):

and that same series.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\sqrt{\frac{x}{1-x^3}}dx}\) \(\color{#000000 }{ \displaystyle \int\frac{\sqrt{x}}{\sqrt{1-x^3}}dx}\) \(\color{#000000 }{ \displaystyle \int\frac{\sqrt{x^4}}{\sqrt{x^3-x^6}}dx}\) will assume x is real, \(\color{#000000 }{ \displaystyle \int\frac{x^2}{\sqrt{x^3-x^6}}dx}\) \(\color{#000000 }{ \displaystyle z=x^3}\) \(\color{#000000 }{ \displaystyle (1/3)dz=x^2~dx}\) \(\color{#000000 }{ \displaystyle \int\frac{1}{3\sqrt{z-z^2}}dz}\) so far..

OpenStudy (xapproachesinfinity):

Solo there a neat solution to second without series

OpenStudy (xapproachesinfinity):

first one don't know thus far, i did think of using euclidean division of x/1-x^3

OpenStudy (solomonzelman):

trying to think of it...

OpenStudy (xapproachesinfinity):

gotta go, good luck in solving them

OpenStudy (solomonzelman):

By parts... \(\color{#000000 }{ \displaystyle \int \ln(\ln x)dx=}\) \(\color{#000000 }{ \displaystyle x\left(\ln(\ln x)\right)-\int x\left(\frac{1}{x\ln x}\right)~dx}\) So, \(\color{#000000 }{ \displaystyle \int \frac{1}{\ln x}+\ln(\ln x)dx=}\) \(\color{#000000 }{ \displaystyle \int \frac{1}{\ln x}dx+\int \ln(\ln x)dx=}\) \(\color{#000000 }{ \displaystyle \int \frac{1}{\ln x}dx+x\left(\ln(\ln x)\right)-\int x\left(\frac{1}{x\ln x}\right)~dx=x\ln(\ln x)}\)

OpenStudy (solomonzelman):

there, @xapproachesinfinity

OpenStudy (solomonzelman):

well, +C, but anyway ...

OpenStudy (solomonzelman):

As far as the second integral goes \(\color{#000000 }{ \displaystyle \int \frac{1}{\sqrt{z-z^2}}dz }\) \(\color{#000000 }{ \displaystyle \int \frac{1}{\sqrt{-z^2+z-\frac{1}{4}+\frac{1}{4}}}dz }\) \(\color{#000000 }{ \displaystyle \int \frac{1}{\sqrt{-\left(z-\frac{1}{2}\right)^2+\frac{1}{4}}}dz }\) w=z-1/2 dw=dz \(\color{#000000 }{ \displaystyle \int \frac{1}{\sqrt{-w^2+\frac{1}{4}}}dw }\) \(\color{#000000 }{ \displaystyle \int \frac{1}{\sqrt{\frac{1}{4}-w^2}}dw }\) \(\color{#000000 }{ \displaystyle w=\frac{1}{2}\sin \theta }\) \(\color{#000000 }{ \displaystyle dw=\frac{1}{2}\cos \theta }\) \(\color{#000000 }{ \displaystyle \int \frac{\frac{1}{2}\cos \theta}{\sqrt{\frac{1}{4}-\frac{1}{4}\sin^2\theta }}d\theta }\) \(\color{#000000 }{ \displaystyle \int d\theta }\)

OpenStudy (xapproachesinfinity):

you go the second excellent!

OpenStudy (xapproachesinfinity):

got*

OpenStudy (solomonzelman):

Thanks;)

OpenStudy (solomonzelman):

didn't put 1/3 there but I hope that is the smallest error I had there.

OpenStudy (solomonzelman):

that's referring to the 1st prob. I like the second problem more tho:) Tnx for posting the problems!

OpenStudy (xapproachesinfinity):

np:)

imqwerty (imqwerty):

:) for 1st question- \(\large t=x^{\frac{3}{2}}\) made life easier<33